TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 8

Created by Main.AntonioVergara on 2007-06-15 - 17:25

Powering Tests Sector 78 - Evening 15.06.2007

Planning for Monday 18.06.2007

  • RQF.A78 and RQD.A78
    • Wait for green light from QPS
    • PIC1 for both circuits
    • Connection of the DC cables to the current leads
    • Connection of a 60A converter to the DC cables in the UA side
    • PCC tests for both circuits

  • RB.78
    • QPS IST. Discharge of the quench heaters, quench loop tests, energy extraction type tests

  • Matching Section
    • Further tests for investigating the triple-quench event to be discussed in MPP meeting on Monday

  • 600A correctors in DFBAO
    • ACR/MEL tests on the damaged cryo-instrumentation cards (RSF2.A78B2, RCS.A78B1)
    • If possible, solution of problem in RCO QPS communication
    • QPS Software upgrade

  • Tests of the 60A CRYO_START/CRYO_MANTAIN signals

Test Report of Friday 15.06.2007

  • Matching Section
    • In the morning RD2.L8 and RQ4.L8 and the 7 correctors ramped to nominal.
    • While starting ramp up, RQ5.L8 quench detector tripped due to the very high acceleration in the converter.
    • After some corrections, RQ5.L8 was also ramped to nominal current. The softaware of the quench detector of this circuit still needs to be upgraded
    • During the afternoon Squeezing tests from LSA were performed

  • 600A correctors in DFBAO
    • Eight out of the nine circuits released by ElQA and ACR went through the PCC tests OK. QPS gathered the data needed to update the detector and released the circuits. PIC2 performed in the afternoon in all of them. Test OK.
    • RCO has a problem with the QPS communication

  • 60A type tests completed OK.

  • RQF.A78 and RQD.A78
    • Switch opening failure issue already solved. Small communication problem to be fixed (world-FIP connection). The two circuits should be released for PCC tests on Monday morning.

  • RB.A78
    • Quench heaters are now charged.

  • Boil-off tests on the DFBAN successfully carried out on Saturday.

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login